OpenStudy (sleepyhead314):

How would I find the average velocity from t=1 to t=3 using the Mean Value Theorem given v(t) = t^2 + 3t - 1 ?

OpenStudy (sleepyhead314):

would it be v(3) - v(1) all over 3-1?

OpenStudy (dan815):

hmmm

OpenStudy (sleepyhead314):

but I have a feeling that that might get me something that is not velocity

OpenStudy (dan815):

no, mvt simply says there exists a point between v(3) and v(1) that is the same as the mean

OpenStudy (dan815):

why do you think that

OpenStudy (sleepyhead314):

m/s /s =/= m/s

OpenStudy (dan815):

what does diving by number hav to do with he units

OpenStudy (dan815):

sec i gotta put stuff downstaira ill help u

OpenStudy (sleepyhead314):

so should I do a(3)-a(1) -------- 3-1 to get average velocity?

OpenStudy (dan815):

that according to your link

OpenStudy (dan815):

is the slope

OpenStudy (dan815):

so its the change in velocity

OpenStudy (sleepyhead314):

my head's going in circles...

OpenStudy (dan815):

ok wait look at htis part

OpenStudy (dan815):

f'(c)=f(b)-f(a) ------- b-a

OpenStudy (dan815):

you se that part

OpenStudy (dan815):

im not sure why tha is true, ill have to see the proof but

OpenStudy (dan815):

this means whereever your mean is, it has that as its tangent line

OpenStudy (dan815):

so lets see where this tangent line belongs on your curve

OpenStudy (sleepyhead314):

if you keep scrolling in the second link, there's a proof

OpenStudy (dan815):

lets draw something here to see what is going on

OpenStudy (dan815):

\[v(t)=t^2+3t-1\\ v(t)=(t+3/2)^2-1-9/4\\ ~~~~~~=(t+3/2)^2-13/4 \]

OpenStudy (sleepyhead314):

*thinks harder* *almost there* ... so should I find the secant slope then set it equal to v'(t) and solve for t which will get me the t value that is when the graph hits the overall velocity average? no I need average velocity value so I plug that t into the original given that should work... I think

OpenStudy (dan815):